Daily Report

CPUC/ANF Environmental Monitoring Activities

Pacific Pipeline Construction Project

Aspen Environmental Group Wednesday, January 27, 1999

The following is a brief summary of the environmental monitoring activities that were completed between January 27, 1999 1:00 and 1:00 on January 28, 1999. This summary is based on reports directly received from the field and has not been edited. If you have any questions please contact (818) 597-3407. Please note the item number and the correction that should be made. This report is intended to serve as a communication document and as an instrument for quality control.

Construction Highlights

01/27/99 (Wednesday)

Segment 1 (Emidio Pump Station to Copco Road): Report for 27 January 1999: Activity was minimal today due to the recent rainy and snowy weather. Crews performed general clean up at Grapevine Pump Station area as well as the Rock Road yard. Dewatering of the surplus pipe strung along Highway 166 was completed today. The successful hydrotest for this pipe was finished yesterday. Welders began cutting the 8" pipe used for dewatering into sections to be hauled off. Minor work occurred at both Emidio and Grapevine Pump Stations. Erosion control was planned at several locations today.

Segment 5 (Lynwood to Long Beach): Report for 1-27-99. Crews pumped out water from the bore pit at the Edison site then centered the vault within the pit. Crews working the Huntway line trenched 60' on the southern section and potholed near the future site of the receiving bore pit. During trenching, an unmarked methane line was encountered. Further , an unmarked water line was damaged adjacent to the ditch line. Meanwhile, crews welded pipe in an effort to complete the tie-in for the 16'' line and lowered the bore machine into the pit within the Tosco refinery. In addition, excavation for the fire hydrants began on the east side of the meter skid. Painting crews worked at the Lynwood station preparing the pipe for bonding while carpenters continued working under the roof section. Segment 4B (Lynwood to El Segundo): Painting and grading crews worked at the El Segundo refinery and a general clean up of the station site continued.

Inspection Summar

In the past 24 hours, 2 inspection event(s) were reported for 2 spread(s). Of these inspections, 0 were non-compliance event(s). Also 0 non-compliance resolution report(s) was resolved.

Non-Compliance Reports (NCR's) Issued

No new non-compliance reports were issued during the last reporting period.

Non-Compliance Resolution Reports (NCRR's) Issued

No new non-compliance reports were issued during the last reporting period.

Summary of Follow-up Items for Today

No follow-up items were reported for today.

Back to Daily Reports

Back to MMCRP Home Page